OpenStudy (anonymous):

Convert to a+bi form. (2cos20 degrees + 2 i sin20 degrees) ^3

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

i have no clue where to start, im so lost.

OpenStudy (mertsj):

I think it may be posted wrong. There should be an "i" in there somewhere.

OpenStudy (anonymous):

yes sorry ive corrected it

OpenStudy (mertsj):

Take the two out.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

\[2(\cos 20+i \sin 20)\]

OpenStudy (anonymous):

2(cos20 degrees + i sin20 degrees) ^3

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Now. Raise the 2 to the third power and multiply the argument by 3

OpenStudy (anonymous):

so 8 times the whole thing in parentheses?

OpenStudy (mertsj):

\[8(\cos 60+i \sin 60)\]

OpenStudy (anonymous):

you just multiply? even though its raised to the power of 3?

OpenStudy (anonymous):

ok and then?

OpenStudy (mertsj):

\[8(\frac{1}{2}+i\frac{\sqrt{3}}{2})=4+4\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

so then multiply the left side again?

OpenStudy (anonymous):

oh whoop sorry never mind

OpenStudy (mertsj):

No. Drop the 2. It should be \[4+4\sqrt{3}i\]

OpenStudy (mertsj):

I forgot. I had already cancelled the 2 into the 8

OpenStudy (anonymous):

okay. thank you. but what if the question was like 4^i ? how would you solve it then

OpenStudy (mertsj):

\[4^i\]

OpenStudy (mertsj):

Like that?

OpenStudy (anonymous):

yes, like that

OpenStudy (anonymous):

log?? why log??

OpenStudy (mertsj):

I think they mean natural log...ln

OpenStudy (anonymous):

but still.. i dont understand why you use natural log?

OpenStudy (anonymous):

actually i understand. what if the question was e^ (iπ)?

OpenStudy (mertsj):

Have you had euler's equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh so cos π +i sinπ ?

OpenStudy (anonymous):

so negative , ok.

OpenStudy (anonymous):

what if it was e^pi/2 i ?

OpenStudy (anonymous):



OpenStudy (mertsj):

Probably treat that like: